OpenStudy (anonymous):

Identify the solution(s)of the system of equations,if any 3x+2y=4, 5x-4y=3

OpenStudy (anonymous):

Take one of the equations, isolate the x or the y, then substitute it into the other equation. x=(3+4y)/5; 3[(3+4y)/5] + 2y = 4. Solve for y, then plug it back into the original equation.

OpenStudy (campbell_st):

why not just double the 1st equation and add it to the 2nd equation. This eliminates y. 6x + 4y = 8 5x - 4y = 3 ----------- 11x = 11 solve for x. When you have x, substitute the value into either equation to find y.

OpenStudy (anonymous):

That works as well, just a different method. You can use substitution or elimination.